The table below shows a book balance sheet for the Wishing Well Motel chain. The company’s long-term debt is secured by its real estate assets, but it also uses short-term bank loans as a permanent source of financing. It pays 10% interest on the bank debt and 8% interest on the secured debt. Wishing Well has 10 million shares of stock outstanding, trading at $88 per share. The expected return on Wishing Well’s common stock is 21%. (Table figures in $ millions.)

Cash and marketable securities $ 160 Bank loan $ 300
Accounts receivable 320 Accounts payable 180
Inventory 50 Current liabilities $ 480
Current assets $ 530
Real estate 2,550 Long-term debt 2,450
Other assets 150 Equity 300
Total $ 3,230 Total $ 3,230


Calculate Wishing Well’s WACC. Assume that the book and market values of Wishing Well’s debt are the same. The marginal tax rate is 21%. (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)
Answer #1

Total market value of Wishing Well = Market Value of long term debt + Market Value of bank loan + Market Value of Equity = [$2,450 + $300 + ($88 * 10)] millions = $3,630 millions

Weight of long term debt = Market Value of long term debt/ Total value = $2,450/$3,630 = 0.674931129476584

Weight of bank loan = Market Value of bank loan/ Total value = $300/$3,630 = 0.082644628099174

Weight of Equity = Market Value of bank loan/ Total value = $880/$3,630 = 0.242424242424242

WACC = 0.674931129476584 * 8% * (1 - 21%) + 0.082644628099174 * 10% * (1 - 21%) + 0.242424242424242 * 21% = 0.04265564738292 + 0.006528925619835 + 0.050909090909091 = 0.100093663911846

Wishing Well’s WACC = 10%